• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Bosted

    Log entry time 23:56:31 on January23,2009

    Entry number 257515

    keyword=Swing Shift SUmmary

    Start shift with negative polarity on BB, HRL -2.35 GeV, half wave plate IN.

    16:40 Change target from reference cell (with 3He 112 psai) to 3He production cell. Checked beam position are o.k. Checked rates are back to standard values. Turned on spin flipper.

    16:49 Start production run 5475 on polarized 3He

    15:15 Started getting target spin direction error messsages. Not really a problem. Xin fixed by re-starting Labview.

    17:45 Suddenly started seeing huge dead time in DAQ. Problem persists for pulser run

    18:10 Asked MCC to turn off beam to diagnose high DT problem more clearly.

    18:20 Decided need to call in Hall to power cycle the trigger supervisor carte.

    18:30 Kalyan went into the Hall on escorted access to try this.

    18:40 Kalyan power cycled all BB crates. Problem seems to have gone away so asked for beam again.

    18:50 Checked everything ok with 2 muA

    19:00 Asked for 14 muA. Start production run 5490. Happex run not good
    for this run, though.

    19:10 Happex DAQ now fixed

    21:45 Noticed some problems with the Happex analysis. Restarted CODA, made some log entries, and paged Bob Micheals, who will look into the problem

    21:50 Lasar spectrum suddenly looks much larger than normal. Paged target expert, Yi.Also noticed oven temperature dropping fast. Yi told us to turn off the lasers, which we did. I also turned off the beam and asked for controlled access. We first checked the triger rates and they looked normal, so apparantly cell is intact. Yi came in about 10 or 15 minutes later, and said the problem is the constant power heater for the oven stopped working.

    22:20 Yi went into the Hall to replace the fuse on the constant power oven heater.

    22:35 The fuse fix worked. Oven temperature is rising. switched to the reference cell and asked for beam permit.

    22:42 Mcc ready to send us beam. Asked for 1 muA. Goal is to get clean events to look at Cherenkov in BB. Adjusted prescale factors and started run 5496

    11:10 Oven seems stable enough. End reference cell run and move target to 3He production target. Checked rates at 2 muA, then went up to 14 muA and started run 5497.


    Good runs this shift:
    Runs comments

    5474 Reference cell with He at 112 psi. 670K

    5475 3He production. Q=0.0138. 250K. Stopeed because of spurious EPICS error regarding target spin direction (not fatal).

    5476. Pol 3He. Q=0.0138

    5490 pol 3He Q=0.0065

    5491 Pol 3He Q=0.053

    5492 pol 3H Q=0.053

    5493 pol 3He Q=0.0034 Target oven stopped working during this run

    5496 Reference cell with 3He 112 psia. Current only 1 muA

    5497 pol 3He1